.A distance of 0.002 m separates two objects of equal mass. If the gravitational


force between them is 0.0104 N, find the mass of each object.

1
Expert's answer
2021-11-10T10:10:56-0500

Explanations & Calculations


"\qquad\qquad
\begin{aligned}
\small F&=\small G\frac{m\times m}{r^2}\\
\small m&=\small r\sqrt{\frac{F}{G}}\\
&=\small 0.002\sqrt{\frac{0.0104N}{6.67\times10^{-11}m^3kg^{-1}s^{-2}}}\\
&=\small 24.97kg\\
&\approx\small 25kg
\end{aligned}"